### v1

@@ 1 Introduction
The concept of spanning-tree congestion was introduced under different names in the late 1990s [1, 30, 28, 13], and in 2003 formalized by Ostrovskii [25], who established some key properties. The problem has been extensively studied since then, and numerous results regarding its graph-theoretic properties and computational complexity have been reported in the literature. Below we review briefly those that are most relevant to our work.
@@ 3 $\mathbb{NP}$-Hardness for Degree-$3$ Graphs
In this section we prove Theorem 1.1. Our proof is via a polynomial-time reduction from an $\mathbb{NP}$-complete version of SAT (defined below), mapping a boolean expression $\phi$ into a graph $G$ and integer $K$, such that $\mathrm{stc}(G)\leq K$ if and only if $\phi$ is satisfiable. $G$ consists of multiple gadget subgraphs, some corresponding to variables and some to clauses, as well as one additional root gadget, with appropriate connections in and between these gadgets. Some gadgets are constructed from smaller sub-gadgets. The most basic gadget is called a double-weight gadget, and it allows us to use edges that are assigned two weight values, with appropriate interpretation. Using double-weighted edges, we construct a more complex flower gadget that will be used as the root gadget and as the gadgets for some clauses. The restriction to degree $3$ makes these constructions quite intricate. A considerably simpler proof for graphs of degree at most $4$ can be found in Appendix A.
@@
An edge $e=(u,v)$ with $\mathop{\mathrm{w}_{1}}(e)=\mathop{\mathrm{w}_{2}}(e)$ can be replaced by $\mathop{\mathrm{w}_{1}}(e)$ edge-disjoint paths from $u$ to $v$ without affecting the congestion or vertex degrees. As shown by Luu and Chrobak [23], this idea can be extended to unequal double weights: in a graph $G$, an edge with weight a:ba\!:\!b (under some mild assumptions) can be replaced by an appropriate double-weight gadget, so that the resulting graph $G^{\prime}$ has the property that $\mathrm{stc}(G)\leq K$ if and only if $\mathrm{stc}(G^{\prime})\leq K$. In their construction, the maximum degree of $G^{\prime}$ becomes large if $b$ is large. Lampis et al. [18] provide a low-degree double-weight gadget, but it is not sufficient for our purpose (because it contains vertices of degree $4$).
The construction of our degree-$3$ double-weight gadget $\mathcal{W}(a,b)$, where $a,b$ represent a double weight a:ba\!:\!b, is given in Appendix B. There, we also prove (see Lemma B.1) that for $a<b$ and $b-a\leq K-2$, $\mathcal{W}(a,b)$ has the following property: if the maximum degree in $G$ is $\Delta\geq 3$ and $G^{\prime}$ is the graph obtained from $G$ by replacing an edge with weight a:ba\!:\!b by a copy of $\mathcal{W}(a,b)$ then (i) the naximum degree in $G^{\prime}$ is also $\Delta$ and (ii) $\mathrm{stc}(G)\leq K$ if and only if $\mathrm{stc}(G^{\prime})\leq K$.
@@
The weights in $G$ will be bounded by a polynomial function of the size of $\phi$; thus, per Lemma B.1 and the construction of the double-weight gadget $\mathcal{W}(a,b)$ in Appendix B, this reduction will be sufficient to establish Theorem 1.1.
@@
  • For each 2P-clause $\gamma$, add one root-clause edge $(\gamma,t_{\mathcal{R}}^{\gamma})$ with weight 1:K11\!:\!K-1. For each 2N-clause $\alpha$, add two root-clause edges $(t_{\alpha}^{\mathcal{R},1},t_{\mathcal{R}}^{\alpha,1})$ and $(t_{\alpha}^{\mathcal{R},2},t_{\mathcal{R}}^{\alpha,2})$, each of weight 1:K11\!:\!K-1.
@@
($\Leftarrow$) Given a spanning tree $T$ for $G$ with $\mathrm{cng}(G,T)\leq K$, our goal is to construct a satisfying assignment for $\phi$. Ideally, $T$ would have a form similar to the spanning tree described in the $\Rightarrow$ direction, but this may not be the case. Nevertheless, $T$ has enough structure to define a variable assignment.
First, we give a suitable definition of connection between clause-gadgets and variable-gadgets that will be used to define our boolean assignment. For a clause $\kappa$ containing variable $x$, a traversal from $\kappa$ to $x$ is a path in $T$ that starts from the center of $\kappa$ and contains exactly two external edges: the first being the clause-variable edge from $\kappa$ to $x$ (the entering edge), and the second being any other external edge adjacent to the $x$-gadget (the exiting edge). We say that $T$ traverses $x$ from $\kappa$ when such a path exists. Intuitively, a traversal from $\kappa$ to $x$ is a natural way to lift the edge $(\kappa,x)$ from $G^{*}$ to a path in $G$. We emphasize that a traversal begins at the centers of flowers, as opposed to terminals or dummy nodes; this becomes relevant later in Claim 3.3.
The following lemma is sufficient to construct an assignment for $\phi$:
@@
Tree $T$ has the following properties: (a) $T$ does not contain any root-clause edges. (b) If $T$ traverses $x$ from a negative clause, then $T$ does not traverse $x$ from any positive clause.
Assuming the lemma above holds, we define our satisfying assignment as follows: for each variable $x$, make it false if $T$ traverses it from a negative clause, otherwise make it true. Lemma 3.2(a) implies that $T$ includes for each clause $\kappa$ a traversal from $\kappa$ to some variable $x$ appearing in it. (In particular, a traversal appears as a prefix of the path in $T$ from $\kappa$’s center to the center of the root.) $x$ is false when $\kappa$ is negative by definition of the assignment, and $x$ is true when $\kappa$ is positive by Lemma 3.2(b). Therefore all clauses are satisfied.
@@
First we prove Lemma 3.2(a). Suppose for contradiction that $T$ contains some root-clause edge $e$ from a two-literal clause $\kappa$. If $\kappa=\alpha$ is a 2N-clause, then $e=(t_{\alpha}^{\mathcal{R},j},t_{\mathcal{R}}^{\alpha,j})$, for $j\in{\left\{{1,2}\right\}}$, and there are two edge-disjoint paths (neither containing $e$) from $t_{\alpha}^{\mathcal{R},j}$ to $t_{\mathcal{R}}^{\alpha,j}$: for each variable $x$ of $\alpha$, walk from $\alpha$ to the $x$-gadget, and then to the root. (By the structure of flowers, discussed earlier, we can ensure that these two paths are indeed edge-disjoint.) This implies that $\mathrm{cng}_{G,T}(e)\geq\mathop{\mathrm{w}_{2}}(e)+2=K-1+2=K+1$, contradicting $\mathrm{cng}_{G}(T)\leq K$. A similar reasoning applies when $\kappa$ is a 2P-clause.
Next we prove Lemma 3.2(b). Let $x$ be a variable and $\alpha,\beta,\gamma$ be the 2N-clause, 3P-clause, and 2P-clause of $x$, respectively. Assume for contradiction that $T$ traverses $x$ from $\alpha$ and at least one positive clause $\beta$ or $\gamma$. We show that this assumption implies that $\mathrm{cng}_{G}(T)>K$. First we prove the following claim.
@@
Continuing the proof of Lemma 3.2(b), let $\pi\in{\left\{{\beta,\gamma}\right\}}$ be a clause satisfying Claim 3.3. By Lemma 3.2(a), the path in $T$ from the center of $\alpha$ to the center of the root must contain some root-variable edge; suppose $e={\left({y^{\mathcal{R}},t_{\mathcal{R}}^{y}}\right)}$ is the first such edge on this path. Clearly, the center of the root lies on the shore $T^{-y^{\mathcal{R}}}_{e}$ of the cut $\partial T_{e}$}, while the center of $\alpha$ lies on the opposite shore $T^{+y^{\mathcal{R}}}_{e}$. On the other hand, the path described in Claim 3.3 does not contain any root-variable edges, implying that $\pi$ and $x^{\mathrm{2N}}$ also lie on shore $T^{+y^{\mathcal{R}}}_{e}$. In conjunction with Observation 3.1, we obtain:
@@ 4.2 Cactus Representation for Graphs with Congestion $K$
Basic $K$-cuts associated with cactus nodes. A key property of $\widetilde{T}$ that we use in our algorithm is that for any basic $K$-cut $\partial X$}, all edges in $\widetilde{T}\cap{\partial X}$ share an endpoint $w$. Lemmas 4.6 and 4.7 below show an even stronger property, namely for any node $b$, for all basic cuts associated with $b$, this common endpoint will be the same; furthermore for Type-1 node $b$, we have $\phi(w)=b$, i.e., we have only this single choice for the common endpoint $w$. See Figure 5 for an illustration.
@@
Next, in the lemma below, we show that $\widetilde{T}$ traverses the cuts represented by $C$ in a very restricted way: for some index $g$, the edges of $\widetilde{T}$ form a length-$(\ell-1)$ path that traverses consecutive half-$K$-cuts from $Z_{g}$ to $Z_{g+\ell-1}=Z_{g-1}$, and the only other edges of $\widetilde{T}$ in the cuts of $C$ are between $w_{g+1}$ and $Z_{g}$ and between $w_{g-2}$ and $Z_{g-1}$. In particular, $\widetilde{T}$ has no edges in the half-cut ${\partial Z_{g-1}}\cap{\partial Z_{g}}$. (See Figure 6.)
@@ Appendix A $\mathbb{NP}$-Hardness for Degree 44
Our construction builds on the ideas from [23, 18]; in particular we use the concept of a graph with double-weighted edges, with weights a:ba\!:\!b that are polynomial in the size of $\phi$. This is permitted by Lemma B.1 and our construction of $\mathcal{W}(a,b)$ in Appendix B. (Alternatively, for this proof we can use the double-weight gadget of degree $4$ constructed in [18].) For brevity, from now on we will refer to double weights simply as “weights”. Similarly, the “degree” of a vertex refers to its weighted degree.
@@
Appendix B The Double-Weight Gadget
In this section we describe the construction of our double-weight gadget $\mathcal{W}(a,b)$, that will be used to replace edges with double weights in our $\mathbb{NP}$-completeness proof in Section 3. The double-weight gadget is constructed from a sub-gadget that we call a bottleneck, that we introduce first.
@@ Appendix B The Double-Weight Gadget
The lemma below, establishing the validity of converting a bottleneck sub-gadget into a double-weight gadget $\mathcal{W}(a,b)$, was shown in [23] for some specific choice of the bottleneck sub-gadget, but it trivially extends to our more general formulation, so we omit its proof here.

### v2

@@ 1 Introduction
The concept of spanning-tree congestion was introduced under different names in the late 1990s [1, 30, 28, 13], and in 2004 formalized by Ostrovskii [25], who established some key properties. The problem has been extensively studied since then, and numerous results regarding its graph-theoretic properties and computational complexity have been reported in the literature. Below we review briefly those that are most relevant to our work.
@@ 3 $\mathbb{NP}$-Hardness for Degree-$3$ Graphs
In this section we prove Theorem 1.1. Our proof is via a polynomial-time reduction from an $\mathbb{NP}$-complete version of SAT (defined below), mapping a boolean expression $\phi$ into a graph $G$ and integer $K$, such that $\mathrm{stc}(G)\leq K$ if and only if $\phi$ is satisfiable. $G$ consists of multiple gadget subgraphs, some corresponding to variables and some to clauses, as well as one additional root gadget, with appropriate connections in and between these gadgets. Some gadgets are constructed from smaller sub-gadgets. The most basic gadget is called a double-weight gadget, and it allows us to use edges that are assigned two weight values, with appropriate interpretation. Using double-weighted edges, we construct a more complex flower gadget that will be used as the root gadget and as the gadgets for some clauses. The restriction to degree $3$ makes these constructions quite intricate. A considerably simpler (through structurally similar) proof for graphs of degree at most $4$ can be found in Appendix A.
@@
An edge $e=(u,v)$ with $\mathop{\mathrm{w}_{1}}(e)=\mathop{\mathrm{w}_{2}}(e)$ can be replaced by $\mathop{\mathrm{w}_{1}}(e)$ edge-disjoint paths from $u$ to $v$ without affecting the congestion or vertex degrees. As shown by Luu and Chrobak [23], this idea can be extended to unequal double weights: in a graph $G$, an edge with weight a:ba\!:\!b (under some mild assumptions) can be replaced by an appropriate unweighted sub-graph called a double-weight gadget, so that the resulting graph $G^{\prime}$ has the property that $\mathrm{stc}(G)\leq K$ if and only if $\mathrm{stc}(G^{\prime})\leq K$. In their construction, the maximum degree of $G^{\prime}$ becomes large if $b$ is large. Lampis et al. [18] provide a low-degree double-weight gadget, but it is not sufficient for our purpose (because it contains vertices of degree $4$).
In Appendix B, we construct a degree $3$ double-weight gadget $\mathcal{W}(a,b),s^{\ast},t^{\ast}$, where $a,b$ represent a double weight a:ba\!:\!b and $s^{\ast},t^{\ast}$ are two designated vertices called “ports”. There we also prove that $\mathcal{W}(a,b)$ has the following properties.
Lemma 3.1.
(ii) $\mathrm{stc}(G)\leq K$ if and only if $\mathrm{stc}(G^{\prime})\leq K$.
@@
The weights in $G$ will be bounded by a polynomial function of the size of $\phi$; thus, per Lemma 3.1 and the construction of the double-weight gadget $\mathcal{W}(a,b)$ in Appendix B, this reduction will be sufficient to establish Theorem 1.1.
@@
  • For each 2P-clause $\gamma$, add one root-clause edge $(\gamma,t_{\mathcal{R}}^{\gamma})$ with weight 1:K11\!:\!K-1. For each 2N-clause $\alpha$, add two root-clause edges $(t_{\alpha}^{\mathcal{R},1},t_{\mathcal{R}}^{\alpha,1})$ and $(t_{\alpha}^{\mathcal{R},2},t_{\mathcal{R}}^{\alpha,2})$, each of weight 1:K11\!:\!K-1. (Note that there are no edges between 3P-clases and the root gadget.)
@@
($\Leftarrow$) Given a spanning tree $T$ for $G$ with $\mathrm{cng}(G,T)\leq K$, our goal is to construct a satisfying assignment for $\phi$. Ideally, $T$ would have a form similar to the spanning tree described in the $\Rightarrow$ direction, but this may not be the case. Nevertheless, $T$ has enough structure, captured in Lemmas 3.3 and 3.4, to define a variable assignment.
@@
Tree $T$ does not contain any root-clause edges.
@@
First we observe that every root-clause edge has weight 1:K11:K-1. Next we observe that for every root-clause edge $(u,v)$, be it an edge between the root and a 2N-clause or 2P-clause, there are (at least) three edge disjoint paths between the vertices $u$ and $v$. Thus, if the edge $(u,v)$ appears in $T$, it has congestion at least K1+2=K+1K-1+2=K+1, contradicting $\mathrm{cng}_{G}(T)\leq K$. We conclude that no root-clause edge appears in $T$. ∎
To state Lemma 3.4, we need a few more definitions. For a clause $\kappa$ containing variable $x$, a traversal from $\kappa$ to $x$ is a path in $T$ that starts from the center of $\kappa$ and contains exactly two external edges: the first being the clause-variable edge from $\kappa$ to $x$ (the entering edge), and the second being any other external edge adjacent to the $x$-gadget (the exiting edge). We say that $T$ traverses $x$ from $\kappa$ when such a path exists. Intuitively, a traversal from $\kappa$ to $x$ is a natural way to lift the edge $(\kappa,x)$ from $G^{*}$ to a path in $G$. We emphasize that a traversal begins at the centers of flowers, as opposed to terminals or dummy nodes; this becomes relevant later in Claim 3.5.
Lemma 3.4.
If $T$ traverses $x$ from a negative clause, then $T$ does not traverse $x$ from any positive clause.
Assuming the lemma above holds, we define our satisfying assignment as follows: for each variable $x$, make it false if $T$ traverses it from a negative clause, otherwise make it true. Lemma 3.3 implies that $T$ includes for each clause $\kappa$ a traversal from $\kappa$ to some variable $x$ appearing in it. (In particular, a traversal appears as a prefix of the path in $T$ from $\kappa$’s center to the center of the root.) $x$ is false when $\kappa$ is negative by definition of the assignment, and $x$ is true when $\kappa$ is positive by Lemma 3.4. Therefore all clauses are satisfied.
Proof.
Let $x$ be a variable and $\alpha,\beta,\gamma$ be the 2N-clause, 3P-clause, and 2P-clause of $x$, respectively. Assume for contradiction that $T$ traverses $x$ from $\alpha$ and at least one positive clause $\beta$ or $\gamma$. We show that this assumption implies that $\mathrm{cng}_{G}(T)>K$. First we prove the following claim.
@@
Continuing the proof of Lemma 3.4, let $\pi\in{\left\{{\beta,\gamma}\right\}}$ be a clause satisfying Claim 3.5. By Lemma 3.3, the path in $T$ from the center of $\alpha$ to the center of the root must contain some root-variable edge; suppose $e={\left({y^{\mathcal{R}},t_{\mathcal{R}}^{y}}\right)}$ is the first such edge on this path. Clearly, the center of the root lies on the shore $T^{-y^{\mathcal{R}}}_{e}$ of the cut $\partial T_{e}$}, while the center of $\alpha$ lies on the opposite shore $T^{+y^{\mathcal{R}}}_{e}$. On the other hand, the path described in Claim 3.5 does not contain any root-variable edges, implying that $\pi$ and $x^{\mathrm{2N}}$ also lie on shore $T^{+y^{\mathcal{R}}}_{e}$. In conjunction with Observation 3.2, we obtain:
@@ 4.2 Cactus Representation for Graphs with Congestion $K$
Basic $K$-cuts associated with cactus nodes. A key property of $\widetilde{T}$ that we use in our algorithm is that for any basic $K$-cut $\partial X$}, all edges in $\widetilde{T}\cap{\partial X}$ share an endpoint $w$. Lemmas 4.6 and 4.7 below show an even stronger property, namely that for all basic cuts associated with any given node $b$, this common endpoint will be the same; furthermore for Type-1 node $b$, we have $\phi(w)=b$, i.e., we have only this single choice for the common endpoint $w$. See Figure 5 for an illustration.
@@
Next, in the lemma below, we show that $\widetilde{T}$ traverses the cuts represented by $C$ in a very restricted way: for some index $g$, the edges of $\widetilde{T}$ form a length-$(\ell-1)$ path that traverses consecutive half-$K$-cuts between sets $Z_{g},Z_{g+1},...,Z_{g+\ell-1}=Z_{g-1}$, and the only other edges of $\widetilde{T}$ in the cuts of $C$ are among those connecting a vertex $w_{g+1}\in Z_{g+1}$ to set $Z_{g}$, and among those connecting a vertex $w_{g-2}\in Z_{g-2}$ to set $Z_{g-1}$. In particular, $\widetilde{T}$ has no edges in the half-cut ${\partial Z_{g-1}}\cap{\partial Z_{g}}$. (See Figure 6.)
@@ Appendix A $\mathbb{NP}$-Hardness for Degree 44
Our construction builds on the ideas from [23, 18]; in particular we use the concept of a graph with double-weighted edges, with weights a:ba\!:\!b that are polynomial in the size of $\phi$. This is permitted by Lemma 3.1 and our construction of $\mathcal{W}(a,b)$ in Appendix B. (Alternatively, for this proof we can use the double-weight gadget of degree $4$ constructed in [18].) For brevity, from now on we will refer to double weights simply as “weights”. Similarly, the “degree” of a vertex refers to its weighted degree.
@@
Appendix B The Double-Weight Gadget – Proof of Lemma 3.1
In this section we prove Lemma 3.1, by describing the construction of our double-weight gadget $\mathcal{W}(a,b)$, that will be used to replace edges with double weights in our $\mathbb{NP}$-completeness proof in Section 3. The double-weight gadget is constructed from a sub-gadget that we call a bottleneck, that we introduce first.
@@
Clearly $\mathcal{W}(a,b)$ satisfies Lemma 3.1(i). That this gadget satisfies Lemma 3.1(ii) was proved by Lampis et al. [18, Lemma 10] when $\mathcal{B}$ is a square-grid, but their argument trivially extends to any bottleneck gadget, so we omit the proof here.